OpenStudy (anonymous):

There were 150 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $405. How many Juniors bought tickets?

OpenStudy (anonymous):

Help again please! @JakeV8 @ktnguyen1

OpenStudy (anonymous):

x is junior y is senior x+y=150 2x+3.5y=450 2x+2y=300 2x+3.5y=450 1.5y=450-300=150 y=100 x=150-100=50 Juniors=50 Seniors=100

OpenStudy (anonymous):

i dont have any of those options in the answers

OpenStudy (anonymous):

wouldnt it be 80 juniors and 70 seniors?

OpenStudy (anonymous):

Oh im sorry 405 i rewrite 450 2x+2y=300 2x+3.5y=405 1.5y=405-300=105 y=70 X=150-70=80 Juniors=80 Seniors=70

OpenStudy (anonymous):

oh ok thx :) I just added a new one please help on that one! really appreciate it!